OpenStudy (anonymous):

Solve the following system of equations. x + y = 14 -x + y = 16 A. (1, 15) B. (1, -15) C. (-1, 15) D. (6, 6

OpenStudy (tester97):

c

OpenStudy (anonymous):

how did you get C ?

OpenStudy (tester97):

Just plug in -1 for x in both equations and plug in 15 for y in both equations

OpenStudy (tester97):

-1+15=14 (yes) -1+15=16 (yes) BAM THERES YOUR ANSWER

OpenStudy (anonymous):

:)